Using a yardstick to measure a metre, you find that one metre equals about 1.09 yards. If you buy some dental floss that is 150.000yard in length, how many metres long is the floss?


1
Expert's answer
2021-02-01T04:03:28-0500

Given , 1 metre = 1.09 yards

So, we can write,

1 yard = 1/1.09 = 0.917 metre

150,000 yards = 0.917 ×150000 = 137.6 metres

Answer = [137.6] metres